### v1

@@ Abstract
journal:
1 Introduction
@@
Let $\mathbb{F}_{q}$ be the finite field with $q$ elements. A subspace $C$ of $\mathbb{F}_{q}^{n}$ refers to an $\mathbb{F}_{q}$-linear code of length $n$. Under the Euclidean inner product, the dual of the code $C$ is the collection of all vectors of $\mathbb{F}_{q}^{n}$ that are orthogonal to $C$ and denoted by $C^{\perp}$. The hull of the code $C$ is defined as $Hull(C)=C\cap C^{\perp}$. In fact, the concept of the hull of an $\mathbb{F}_{q}$-linear code was introduced in 1990 by Assmus and Key Assmus to investigate certain properties of finite projective planes. Recall that two codes $C$ and $D$ are permutation-equivalent if $C$ can be derived from $D$ by applying a suitable coordinate permutation. Determining whether two linear codes are permutation-equivalent is a significant problem in coding theory. Recently, hulls have attracted intense research interest due to their various applications. In particular, the dimension of the hull directly influences the complexity of algorithms that check the permutation equivalence or determine the automorphism group of a linear code Leon ; Send1997 ; Send2000 . Additionally, hulls can be utilized to construct quantum error-correcting codes Dougherty ; Guenda . These facts highlight the necessity of studying hull codes. Over the years, researchers have investigated the hulls of various codes over finite fields or finite rings with unity. However, there is no work available in the literature on the hull of codes over non-unital rings. Thus, the study of hulls over non-unital rings is a significant problem, and this is a first attempt towards this direction.
@@
Recall that an $\mathbb{F}_{q}$-linear code $C$ is referred to as a linear complementary dual (LCD), self-orthogonal (SO) and self-dual (SD) code if it satisfies $\cap C^{\perp}=\{0\}, CC$$\subseteq C^{\perp}$ and $C=C^{\perp}$, respectively. We see that, if $Hull(C)$ is trivial, it is LCD, if $Hull(C)=C$, it is SO, and if $Hull(C)=C=C^{\perp}$, it is SD. Thus, the hull code can be seen as a basic generalization of LCD, SO and SD codes. The concept of LCD codes was introduced by Massey in Massey92 , where he also proved the existence of asymptotically good LCD codes. In recent years, these codes have attracted considerable interest because of their use in countering Side-Channel Attacks Carl16 . Using the concept of hull dimension spectra, Sendrier proved in 2004 that LCD codes attain the Gilbert-Varshamov bound Send04 . Subsequently, Islam et al. Islam2022 demonstrated a method to construct both quantum and LCD codes leveraging cyclic codes over non-chain rings. Further, double circulant LCD and SD codes were investigated in Prakash . On the other hand, research on SO and SD codes has seen significant growth due to their use in different areas such as quantum stabilizer codes Calderbank ; Steane , modular forms Minjia , lattice theory Bannai ; Conway ; Harada , and combinatorial $t$-design theory Bachoc .
@@
Recently, there has been a significant shift towards considering non-unital rings to study different types of codes. In 2022, Shi et al. Shi21 were the first to consider a non-unitary ring with $4$ elements to investigate LCD codes. Following this, Kim and Roe Kim2022a investigated quasi-self-dual (QSD) codes over a non-unitary ring with 4 elements. Note that a QSD code $C$ of length $n$ over a finite ring $R$ is an SO code with $|C|=|R|^{n/2}$. Subsequently, authors in Kushwaha1 extended QSD and LCD codes over a non-unitary ring with $9$ elements. Meanwhile, Deb et al. Deb worked on classifying certain codes over a non-unitary ring of order $4$. This was followed by the work of Kushwaha et al. Kushwaha2 , who classified MDS and almost MDS LCD and SD codes over a non-unital ring $E_{p}$. On the other hand, the build-up construction method is a powerful technique to construct codes with a larger length from codes with a smaller length. This technique was utilized in Alah22 and Kim2022a for the classification of Type IV and QSD codes over non-unital rings for some smaller lengths. Further, binary linear codes with various hull dimensions were classified in Li using build-up construction techniques. In 2020, Liu and Pan Liu studied the hull-variation problem for $\mathbb{F}_{q}$-linear codes, while LCD and $\mathbb{F}_{q}$-linear codes with hull-dimension $1$ were investigated in Wang . Motivated by these works, we study the hulls of free codes over the non-unital ring $E=\langle\kappa,\tau\mid 2\kappa=2\tau=0,~\kappa^{2}=\kappa,~\tau^{2}=\tau,~\kappa\tau=\kappa,~\tau\kappa=\tau\rangle$. To achieve this, we explore the various hulls of the $E$-linear codes and identify their residue and torsion codes. Then, we produce the form of the generator matrix of the hull code of a free $E$-linear code and also determine its hull-rank. Subsequently, we present four build-up construction methods that utilize free $E$-linear codes with a smaller length and hull-rank to construct free $E$-linear codes with a larger length and hull-rank. We also give some supporting examples of codes constructed by using these build-up construction methods. In addition, we investigate the permutation equivalence of $E$-linear codes and discuss the hull-variation problem for free $E$-linear codes. We conclude our work by classifying optimal free $E$-linear codes for lengths up to $8$.
@@
2 Hulls over the ring $E$
In Fine93 , Fine has classified rings of order $4$ up to isomorphism. There are only two noncommutative non-unitay rings with $4$ elements, namely, the ring $E=\langle\kappa,\tau\mid 2\kappa=2\tau=0,~\kappa^{2}=\kappa,~\tau^{2}=\tau,~\kappa\tau=\kappa,~\tau\kappa=\tau\rangle$ and its opposite ring. This paper considers the ring $E$ as a code alphabet to study the hull of free linear codes over $E$. One can derive similar results over the opposite ring. The characteristic of the ring $E$ is $2$ and it has $4$ elements $\{~i\kappa+j\tau~|~0\leq i,j<2~\}$. We denote $\kappa+\tau=\zeta$ and give the multiplication table of the ring $E$ in Table 1 to characterize its structure.
Table 1: Multiplication table of $E$.
@@ 2 Hulls over the ring $E$
Definition 1(Linear code)
@@
Definition 2(Minimum distance)
@@
Definition 3(Left hull, right hull and hull codes)
@@
Definition 4(Generating set, generator matrix and parity-check matrix)
@@
Definition 5(Free code)
@@
Definition 6(Rank of a free code)
@@
We now state a few basic results derived from Lemma $1$, Theorem $9$ and Theorem $10$ of Alah23 .
Theorem 2.1
@@
Proof
Freeness of $C^{\perp_{L}}$ and non-freeness of $C^{\perp_{R}}$ follow the Lemma $5$ of Shi21 . To prove the necessity of the second result, i.e., for the free $E$-linear code $C$, the result (4) of Theorem 2.1 implies that
@@
Therefore, $C^{\perp}$ is free. On the other hand, assume $C^{\perp}$ is a free $E$-linear code. Since $(C^{\perp})_{Res}=(C^{\perp})_{Tor}$, again (4) of Theorem 2.1 implies that $(C_{Res})^{\perp}=(C_{Tor})^{\perp}$. Thus, $C_{Res}=C_{Tor}$, and so $C$ is free.
@@
∎An important consequence of the above theorem is the following corollary.
Corollary 1
@@
Thus, the left hull and two-sided hull codes of the free $E$-code $C$ are also free.
@@
Thus, the left hull and two-sided hull codes of the free $E$-code $C$ are equal.
@@
Proof
Since $C$ is a free $E$-code, $Hull(C)$ is also free by Theorem 2.6. Next, by Corollary 1, we have $(Hull(C))_{Res}=Hull(C_{Res})$. Moreover, from Theorem $1$ of Kushwaha2 , if $C$ is a free $E$-linear code and $G_{1}$ is a generator matrix of $C_{Res}$, then $\kappa G_{1}$ is a generator matrix of $C$. Therefore, if the hull code of $C_{Res}$ has a generator matrix $G$, then $\kappa G$ is a generator matrix of $Hull(C)$.
@@
Proof
Follows the definition of the rank of a free $E$-linear code and Theorem 2.8.
3 Build-Up Constructions
@@
Now, we recall Proposition $1$ of Li18 for our further investigations.
Proposition 1
@@
    $rank(G_{1}G_{1}^{T})=rank(G^{\prime}_{1}(G^{\prime}_{1})^{T})=rank(M)$ $=rank(G_{2}G_{2}^{T})=rank(GG^{T})=k-l.$
@@
    $rank(G_{1}G_{1}^{T})=rank(G^{\prime}_{1}(G^{\prime}_{1})^{T})=rank(M)=rank(G_{2}G_{2}^{T})-1=rank(GG^{T})-1=k-l-1.$
@@
    $rank(G_{1}G_{1}^{T})=rank(G^{\prime}_{1}(G^{\prime}_{1})^{T})=rank(M)=rank(G_{2}G_{2}^{T})+1=rank(GG^{T})+1=k-l+1.$
@@ 4 Classification
Definition 7(Permutation-equivalent codes)
@@
The following result from Alah23 investigates the permutation equivalence of two free $E$-linear codes.
Theorem 4.1
( Alah23 , Theorem $16$) Two free $E$-linear codes $C$ and $D$ are permutation-equivalent if and only if $C_{Res}$ and $D_{Res}$ are permutation-equivalent.
@@
Definition 8($l_{i}$-optimal free code)
@@
Here, we classify optimal free $E$-linear codes with a fixed hull-rank. Since LCD $E$-codes are free $E$-linear codes whose hull has rank 0, and their classification is already provided in Kushwaha2 . Therefore, we classify only the free $E$-linear codes with non-zero hull-ranks. We use Theorem $1$ of Kushwaha2 to get all the free $E$-linear codes, and then use Theorem 4.1 to find all the permutation-inequivalent free linear codes over $E$. Further, Theorem 2.9 helps to find their hull-ranks. Later, we use the Definition 8 to determine their optimality and list them in Tables 3, 3, 5, 5, 7, 7, 9, 9 and 10. In the tables, $I_{m}$ is the identity matrix of order $m$. Here, all computations have been carried out using MAGMA software Magma .
Table 2: $l_{1}$-optimal free codes (Part I).
@@
Table 3: $l_{1}$-optimal free codes (Part II).
Generator Matrix [$n,k,d$]
@@
Table 4: $l_{1}$-optimal free codes (Part III).
Generator Matrix $[n,k,d]$
@@
Table 5: $l_{1}$-optimal free codes (Part IV).
Generator Matrix [$n,k,d$]
@@
Table 6: $l_{1}$-optimal free codes (Part V).
@@
Table 7: $l_{2}$-optimal free codes (Part I).
@@
Table 8: $l_{2}$-optimal free codes (Part II).
Generator Matrix $[n,k,d]$
@@
Table 9: $l_{3}$-optimal free codes.
@@
Table 10: $l_{4}$-optimal free codes.
@@
5 Conclusion
In this paper, we have studied hulls of free linear codes over a non-unital ring $E$. Initially, we have focused on residue and torsion codes of various hulls and found the generator matrix of the hull of a free $E$-linear code. Then, four build-up construction methods have been given for constructing free $E$-linear codes with a larger length and hull-rank from free $E$-linear codes with a smaller length and hull-rank. Some examples of codes constructed by these build-up construction methods are also given. Later, we studied the permutation equivalence of two free $E$-linear codes and then discussed the hull-variation problem. Finally, we have classified optimal free $E$-linear codes under permutation equivalence for lengths up to $8$. We would also like to point out that extending this work to other non-unital rings, as appeared in the classification of Fine Fine93 , can be an important direction for future research.

### v2

@@ Abstract.
Key words and phrases:
Non-unital rings, Hull, Build-up construction, Hull-variation problem
1991 Mathematics Subject Classification:
94B05, 16L30
* Corresponding author
1. Introduction
@@
Let $\mathbb{F}_{q}$ be the finite field with $q$ elements. A subspace $C$ of $\mathbb{F}_{q}^{n}$ refers to an $\mathbb{F}_{q}$-linear code of length $n$. Under the Euclidean inner product, the dual of the code $C$ is the collection of all vectors of $\mathbb{F}_{q}^{n}$ that are orthogonal to $C$ and denoted by $C^{\perp}$. The hull of the code $C$ is defined as $Hull(C)=C\cap C^{\perp}$. In fact, the concept of the hull of an $\mathbb{F}_{q}$-linear code was introduced in 1990 by Assmus and Key [4] to investigate certain properties of finite projective planes. Recall that two codes $C$ and $D$ are permutation-equivalent if $C$ can be derived from $D$ by applying a suitable coordinate permutation. Determining whether two linear codes are permutation-equivalent is a significant problem in coding theory. Recently, hulls have attracted intense research interest due to their various applications. In particular, the dimension of the hull directly influences the complexity of algorithms that check the permutation equivalence or determine the automorphism group of a linear code [20, 27, 28]. Additionally, hulls can be utilized to construct quantum error-correcting codes [12, 22]. These facts highlight the necessity of studying hull codes. Over the years, researchers have investigated the hulls of various codes over finite fields or finite rings with unity. However, there is no work available in the literature on the hull of codes over non-unital rings. Thus, the study of hulls over non-unital rings is a significant problem, and this is a first attempt towards this direction.
@@
Recall that an $\mathbb{F}_{q}$-linear code $C$ is referred to as a linear complementary dual (LCD), self-orthogonal (SO) and self-dual (SD) code if it satisfies $\cap C^{\perp}=\{0\}, CC$$\subseteq C^{\perp}$ and $C=C^{\perp}$, respectively. We see that, if $Hull(C)$ is trivial, it is LCD, if $Hull(C)=C$, it is SO, and if $Hull(C)=C=C^{\perp}$, it is SD. Thus, the hull code can be seen as a basic generalization of LCD, SO and SD codes. The concept of LCD codes was introduced by Massey in [24], where he also proved the existence of asymptotically good LCD codes. In recent years, these codes have attracted considerable interest because of their use in countering Side-Channel Attacks [8]. Additionally, LCD codes have applications in secret sharing schemes [33]. Using the concept of hull dimension spectra, Sendrier proved in 2004 that LCD codes attain the Gilbert-Varshamov bound [29]. Subsequently, Islam et al. [15] demonstrated a method to construct both quantum and LCD codes leveraging cyclic codes over non-chain rings. Further, double circulant LCD and SD codes were investigated in [25]. On the other hand, research on SO and SD codes has seen significant growth due to their use in different areas such as quantum stabilizer codes [10, 31], modular forms [30], lattice theory [7, 9, 14], and combinatorial $t$-design theory [6].
@@
Recently, there has been a significant shift towards considering non-unital rings to study different types of codes. In 2021, Alahmadi et al. [1] studied Quasi Type IV codes over a non-unital ring. Then, in 2022, Shi et al. [26] were the first to consider a non-unitary ring with $4$ elements to investigate LCD codes. Following these, Kim and Roe [16] investigated quasi-self-dual (QSD) codes over a non-unitary ring with 4 elements. Note that a QSD code $C$ of length $n$ over a finite ring $R$ is an SO code with $|C|=|R|^{n/2}$. Subsequently, authors in [17] extended QSD and LCD codes over a non-unitary ring with $9$ elements. Meanwhile, Deb et al. [11] worked on classifying certain codes over a non-unitary ring of order $4$. This was followed by the work of Kushwaha et al. [18], who classified MDS and almost MDS LCD and SD codes over a non-unital ring $E_{p}$. On the other hand, the build-up construction method is a powerful technique to construct codes with a larger length from codes with a smaller length. This technique was utilized in [2] and [16] for the classification of Type IV and QSD codes over non-unital rings for some smaller lengths. Further, binary linear codes with various hull dimensions were classified in [23] using build-up construction techniques. In 2020, Liu and Pan [21] studied the hull-variation problem for $\mathbb{F}_{q}$-linear codes, while LCD and $\mathbb{F}_{q}$-linear codes with hull-dimension $1$ were investigated in [32]. Motivated by these works, we study the hulls of free codes over the non-unital ring $E=\langle\kappa,\tau\mid 2\kappa=2\tau=0,~\kappa^{2}=\kappa,~\tau^{2}=\tau,~\kappa\tau=\kappa,~\tau\kappa=\tau\rangle$. To achieve this, we explore the various hulls of the $E$-linear codes and identify their residue and torsion codes. Then, we produce the form of the generator matrix of the hull code of a free $E$-linear code and also determine its hull-rank. Subsequently, we present four build-up construction methods that utilize free $E$-linear codes with a smaller length and hull-rank to construct free $E$-linear codes with a larger length and hull-rank. We also give some supporting examples of codes constructed by using these build-up construction methods. In addition, we investigate the permutation equivalence of $E$-linear codes and discuss the hull-variation problem for free $E$-linear codes. We conclude our work by classifying optimal free $E$-linear codes for lengths up to $8$.
@@
2. Hulls over the ring $E$
In [13], Fine has classified rings of order $4$ up to isomorphism. There are only two noncommutative non-unitay rings with $4$ elements, namely, the ring $E=\langle\kappa,\tau\mid 2\kappa=2\tau=0,~\kappa^{2}=\kappa,~\tau^{2}=\tau,~\kappa\tau=\kappa,~\tau\kappa=\tau\rangle$ and its opposite ring. This paper considers the ring $E$ as a code alphabet to study the hull of free linear codes over $E$. One can derive similar results over the opposite ring. The characteristic of the ring $E$ is $2$ and it has $4$ elements $\{~i\kappa+j\tau~|~0\leq i,j<2~\}$. We denote $\kappa+\tau=\zeta$ and give the multiplication table of the ring $E$ in Table 1 to characterize its structure.
Table 1. Multiplication table of $E$.
@@ 2. Hulls over the ring $E$
Definition 1 (Linear code).
@@
Definition 2 (Minimum distance).
@@
Definition 3 (Left hull, right hull and hull codes).
@@
Definition 4 (Generating set, generator matrix and parity-check matrix).
@@
Definition 5 (Free code).
@@
Definition 6 (Rank of a free code).
@@
We now state a few basic results derived from Lemma $1$, Theorem $9$ and Theorem $10$ of [3].
Theorem 2.2.
@@
Proof.
Freeness of $C^{\perp_{L}}$ and non-freeness of $C^{\perp_{R}}$ follow the Lemma $5$ of [26]. To prove the necessity of the second result, i.e., for the free $E$-linear code $C$, the result (4) of Theorem 2.2 implies that
@@
Therefore, $C^{\perp}$ is free. On the other hand, assume $C^{\perp}$ is a free $E$-linear code. Since $(C^{\perp})_{Res}=(C^{\perp})_{Tor}$, again (4) of Theorem 2.2 implies that $(C_{Res})^{\perp}=(C_{Tor})^{\perp}$. Thus, $C_{Res}=C_{Tor}$, and so $C$ is free. ∎
@@
An important consequence of the above theorem is the following corollary.
Corollary 1.
@@
Thus, the left hull and two-sided hull codes of the free $E$-code $C$ are also free. ∎
@@
Thus, the left hull and two-sided hull codes of the free $E$-code $C$ are equal. ∎
@@
Proof.
Since $C$ is a free $E$-code, $Hull(C)$ is also free by Theorem 2.7. Next, by Corollary 1, we have $(Hull(C))_{Res}=Hull(C_{Res})$. Moreover, from Theorem $1$ of [18], if $C$ is a free $E$-linear code and $G_{1}$ is a generator matrix of $C_{Res}$, then $\kappa G_{1}$ is a generator matrix of $C$. Therefore, if the hull code of $C_{Res}$ has a generator matrix $G$, then $\kappa G$ is a generator matrix of $Hull(C)$. ∎
@@
Proof.
Follows the definition of the rank of a free $E$-linear code and Theorem 2.9. ∎
3. Build-Up Constructions
@@
Now, we recall Proposition $1$ of [19] for our further investigations.
Proposition 1.
@@
    $rank(G_{1}G_{1}^{T})$ $=rank(G^{\prime}_{1}(G^{\prime}_{1})^{T})$
    $=rank(M)$
    $=rank(G_{2}G_{2}^{T})$
    $=rank(GG^{T})$
    $=k-l.$
@@
    $rank(G_{1}G_{1}^{T})$ $=rank(G^{\prime}_{1}(G^{\prime}_{1})^{T})$
    $=rank(M)=rank(G_{2}G_{2}^{T})-1$
    $=rank(GG^{T})-1$
    $=k-l-1.$
@@
    $rank(G_{1}G_{1}^{T})$ $=rank(G^{\prime}_{1}(G^{\prime}_{1})^{T})$
    $=rank(M)$
    $=rank(G_{2}G_{2}^{T})+1$
    $=rank(GG^{T})+1$
    $=k-l+1.$
@@ 4. Classification
Definition 7 (Permutation-equivalent codes).
@@
The following result from [3] investigates the permutation equivalence of two free $E$-linear codes.
Theorem 4.1.
( [3], Theorem $16$) Two free $E$-linear codes $C$ and $D$ are permutation-equivalent if and only if $C_{Res}$ and $D_{Res}$ are permutation-equivalent.
@@
Definition 8 ($l_{i}$-optimal free code).
@@
Here, we classify optimal free $E$-linear codes with a fixed hull-rank. Since LCD $E$-codes are free $E$-linear codes whose hull has rank 0, and their classification is already provided in [18]. Therefore, we classify only the free $E$-linear codes with non-zero hull-ranks. We rely on [23] for binary linear codes with a fixed hull-dimension and use Theorem $1$ of [18] to get all the free $E$-linear codes, and then use Theorem 4.1 to find all the permutation-inequivalent free linear codes over $E$. Further, Theorem 2.10 helps to find their hull-ranks. Later, we use the Definition 8 to determine their optimality and list them in Tables 3, 3, 5, 5, 7, 7, 9, 9, 12, 12, 12, 14 and 14. In the tables, $I_{m}$ is the identity matrix of order $m$. Here, all computations have been carried out using MAGMA [5].
Table 2. $l_{1}$-optimal free codes (Part I).
@@
Table 3. $l_{1}$-optimal free codes (Part II).
Generator Matrix [$n,k,d$]
@@
Table 4. $l_{1}$-optimal free codes (Part III).
Generator Matrix $[n,k,d]$
@@
Table 5. $l_{1}$-optimal free codes (Part IV).
Generator Matrix [$n,k,d$]
@@
Table 6. $l_{1}$-optimal free codes (Part V).
Generator Matrix $[n,k,d]$
@@
Table 7. $l_{1}$-optimal free codes (Part VI).
Generator Matrix $[n,k,d]$
@@
Table 8. $l_{1}$-optimal free codes (Part VII).
@@
Table 9. $l_{2}$-optimal free codes (Part I).
@@
Table 10. $l_{2}$-optimal free codes (Part II).
Generator Matrix [$n,k,d$]
@@
Table 11. $l_{2}$-optimal free codes (Part III).
Generator Matrix [$n,k,d$]
@@
Table 12. $l_{3}$-optimal free codes (Part I).
@@
Table 13. $l_{3}$-optimal free codes (Part II).
Generator Matrix [$n,k,d$]
@@
Table 14. $l_{4}$-optimal free codes.
@@
5. Conclusion
In this paper, we have studied hulls of free linear codes over a non-unital ring $E$. Initially, we have focused on residue and torsion codes of various hulls and found the generator matrix of the hull of a free $E$-linear code. Then, four build-up construction methods have been given for constructing free $E$-linear codes with a larger length and hull-rank from free $E$-linear codes with a smaller length and hull-rank. Some examples of codes constructed by these build-up construction methods are also given. Later, we studied the permutation equivalence of two free $E$-linear codes and then discussed the hull-variation problem. Finally, we have classified optimal free $E$-linear codes under permutation equivalence for lengths up to $8$. We would also like to point out that extending this work to other non-unital rings, as appeared in the classification of Fine [13], can be an important direction for future research.